# A Review on Deep Learning Autoencoder in the Design of Next-Generation Communication Systems

### v1

@@ X-A1 Transceiver Design
X-A2 Channel Coding and Decoding
The importance of AEs for channel coding and decoding stems from their potential to redefine communication system design through end-to-end optimisation. Traditional coding approaches often rely on separate stages for modulation and channel coding, optimised under conventional assumptions. However, such designs struggle to adapt to complex or non-conventional channel conditions, leading to suboptimal performance. AEs, using deep learning, integrate channel coding and decoding within a unified framework. By learning directly from data, they overcome limitations of traditional methods, achieving near-optimal performance in challenging environments like non-AWGN channels, mismatched conditions, or inter-symbol interference. Studies such as [75] and [76] demonstrate how AEs improve robustness, reliability, and efficiency, paving the way for advanced adaptive communication systems.
Turbo AEs and Hybrid Designs
Jiang et al. [75] proposed the Turbo AE (TurboAE), which attained state-of-the-art performance in conventional AWGN channels for moderate block lengths and surpassed traditional codes such as LDPC and Turbo codes in non-standard settings like additive T-distribution noise, thereby demonstrating the potential of deep learning techniques to automate the design of channel coding design. Based on this, Jiang et al. [77] introduced TurboAE-MOD, which combines channel coding and modulation within an end-to-end framework. This approach exhibited enhanced reliability and robustness, especially under non-Gaussian channel conditions, when compared with conventional modulation techniques.
Channel AEs and Model-Free Approaches
Zou et al. [76] provided a detailed overview of channel AEs, highlighting the distinctions between model-free approaches (e.g., GAN-CAE) and model-assumed methods (e.g., AWGN-CAE), both of which demonstrated competitive BLER performance, although model-free methods required extended training due to their reliance on surrogate models. Complementing this, Xu et al. [78] evaluated end-to-end AEs in the AWGN and Rayleigh channels, achieving performance comparable to Hamming codes, while exhibiting improved robustness in practical scenarios characterised by channel mismatches, notably within urban and suburban environments.
AEs for ISI Channels and Blind Equalisation
Zhang et al. [79] introduced a Bi-GRU-based AE tailored for ISI channels, offering flexible coding rates and outperforming convolutional and LDPC codes in low SNR regimes. Nevertheless it requires further enhancements at high SNR. Similarly, Caciularu et al. [80] introduced VAE Equalisers specifically for the purpose of blind equalisation, which resulted in considerable enhancements in SER and exhibited strong performance in nonlinear ISI channels, thus outperforming established methodologies such as Turbo EM.
Innovations in Joint Source-Channel Coding (JSCC)
Saidutta et al. [81] developed a VAE-based JSCC scheme, which achieved near-optimal SDR for Gaussian sources while maintaining robustness across different channel conditions. The approach also demonstrated superior performance on image datasets compared to traditional methods.
Specialized AE Implementations
Xu et al. [82] introduced a multi-carrier turbo-style AE for OFDM systems, which delivered significant performance improvements over conventional OFDM and OFDM-IM while allowing seamless integration with existing systems. Rajapaksha et al. [83] developed a low-complexity AE that outperformed convolutional coding in the low to moderate SNR ranges, which achieved notable parameter reduction. Furthermore, Balevi et al. [84] proposed a system combining turbo codes with AE for one-bit quantised receivers, which achieved superior BER and enhanced computational efficiency compared to traditional coding methods.
Advancements in Capacity-Approaching Codes
Letizia et al. [85] proposed an AE loss function that incorporates mutual information to design capacity-approaching codes. The study highlighted advances in BLER and mutual information optimisation, although specific performance values depend on channel setup.
Robust Transmission of Latent Representations
Khan et al. [86] combined basic and denoising AE to enhance robustness in transmitting latent image representations under noise and quantisation. This approach improved classification accuracy and reconstruction quality under challenging conditions.
TABLE VII: Comparison of AE-Based Models in Channel Coding and Decoding
[75] Proposes Turbo AE with state-of-the-art performance for conventional and non-conventional settings TurboAE conventional and Non-AWGN Channels BER $<10^{-5}$ at high SNR; Outperforms benchmarks in reliability for non-conventional channels Slightly worse performance than LDPC and Polar codes at high SNR
[76] Discusses model-free and model-assumed channel AEs, with extensive performance comparison Channel AEs (GAN-CAE, RL-CAE) AWGN, Rayleigh Channels Comparable BLER under AWGN and Rayleigh channels; Model-free methods robust to complex channels Model-free methods require more training epochs
[79] Proposes a flexible AE for coding with arbitrary rates Bi-GRU-based AE ISI Channels, AWGN Outperforms convolutional and LDPC codes in low SNR Performance degrades in high SNR
[80] Introduces a VAE-based approach for blind equalisation and decoding VAE Linear/Nonlinear ISI Channels BER $10^{-3}$ at SNR 6 dB; Robust against nonlinear channels Computational complexity increases with channel impulse length
@@ X-A2 Channel Coding and Decoding
[82] Introduces a flexible MC-AE for integration with any channel coding Turbo-style Multi-Carrier AE OFDM Systems Achieves frequency diversity gain and surpasses OFDM-IM performance High computational cost due to multi-stage learning
[81] Proposes a JSCC approach using VAE for various source distributions VAE with Mixture-of-Experts Gaussian and Laplace Sources; AWGN Channels SDR within 0.96 dB of Shannon limit for Gaussian sources Performance varies with source-channel mismatch
[85] Incorporates mutual information in loss for capacity-approaching codes Capacity-Approaching AE Unknown Channels Significant advancements in BLER and mutual information Results rely on channel assumptions
[83] Proposes a low-complexity AE for coded systems Low-Complexity AE AWGN Channels Better BER than 16-QAM in low SNR; MER $<10^{-4}$ at 10 dB Performance lags in high SNR
[77] Combines TurboAE and feed-forward NN for modulation coding TurboAE-MOD AWGN, Additive T-Distributed Noise Comparable to Turbo codes at moderate block lengths; Superior in non-AWGN channels Degrades in high SNR
[86] Improves robustness against noise and quantisation for image transmission Basic AE + Denoising AE Latent Representations of Images Maximum CA and minimum MSE with latent vector size 20, 4 bits High sensitivity to quantisation
[84] Combines Turbo code and AE for one-bit quantized receivers ConvAE with Turbo Code AWGN Channels Lowest BER for 4-bit transmission; Processes batch in 12.2 µs Training requires careful tuning
[78] Evaluates AE performance across channel types End-to-End AE AWGN, Rayleigh, Urban/Suburban Channels Comparable BLER to Hamming codes; Robust under channel mismatch Slightly worse in high-speed channels
X-A3 Modulation and Demodulation
AEs have emerged as transformative tools in wireless communication, particularly for tasks such as modulation and demodulation. These deep learning models replace traditional transmitter and receiver modules with an encoder-decoder architecture, enabling end-to-end optimisation of communication systems. Unlike conventional methods that optimise modulation and coding separately, AEs jointly learn these processes, allowing them to dynamically adapt to channel impairments and environmental variations. This holistic approach not only improves the efficiency and resilience of the system, but also lays the foundation for robust communication systems capable of addressing the complexities of modern wireless networks.
Signal Transmission and Encoding
Building on this foundational concept, Park et al. [87] demonstrated the potential of AEs for signal transmission in sonobuoy systems. Their approach encoded signals into low-dimensional latent vectors, which reduced data volume by 100 times while enhancing security. Reconstruction errors were kept below 4%, with effective noise reduction achieved using a denoising AE, resulting in a MSE of 0.000238 with the AE model.
Spatial Modulation (SM)
Shrestha et al. [88] tackled challenges in SM by embedding antenna index signatures within AE frameworks, which significantly improved signal robustness under high antenna correlation and achieved BLER and power efficiency gains of 18--24 dB in Rician channel setups.
SNR Improvement
Further advancing the field, Duan et al. [89] developed a dual-residual denoising AE with channel attention (DRdA-CA), which achieved substantial improvements in SNR, such as a gain of 8.38 dB for 16-QAM, and reduced BER by up to 63.5% . Zhu et al. [90] complemented this by presenting a stacked denoising AE for rapid and accurate modulation classification, outperforming traditional methods, especially in noisy environments. Expanding on SDAE applications, Bogale et al. [91] applied this technique to FSO communications for signal denoising and used a deep belief network (DBN) for wavelength demodulation, which achieved notable SNR improvements of +22 dB and high wavelength demodulation accuracy of 0.06 nm.
Modulation Classification
In terms of modulation classification, Ali et al. [92] introduced an AE with non negativity constraints, which achieved 100% classification accuracy at high SNR (10--15 dB) and demonstrated robustness under fading conditions. This approach outperformed traditional sparse AEs in both reconstruction error and classification accuracy. Based on this, Ke et al. [93] designed a LSTM-based denoising AE for radio technology and modulation classification, achieving a high accuracy of 96.56% at high SNR in the RadioML2018.01A dataset. Their model also demonstrated computational efficiency, operating effectively on low-cost platforms such as the Raspberry Pi 4.
Demodulation and Signal Reconstruction
Finally, Zhang et al. [94] introduced hybrid feature-modulated AEs to improve the demodulation of high-frequency information in channel-modulated polarisation imaging. Their method achieved notable improvements in metrics such as PSNR (e.g. $+$ 180% for S3) and SSIM scores, which proved to be effective for simulated and real-world images. Morocho et al. [39] further showcased the versatility of AEs by developing an end-to-end communication system that replaced traditional transmitter and receiver tasks. Their approach optimised system operations and was effectively adapted to channel impairments, highlighting the broad applicability of AE frameworks in communication systems.
TABLE VIII: Comparison of AE-Based Studies for Modulation and Demodulation
Ref. Main Contribution Model(s) Used Dataset/Data Type Best Performance Value Limitation
[39] Joint optimisation of transmitter and receiver tasks for adaptable communication systems. End-to-End AE Single-antenna systems with channel impairments Demonstrated flexibility in handling channel impairments. Lack of quantitative performance metrics for comparison with conventional systems.
[87] Signal transmission with reduced data volume (100x reduction) and enhanced security using low-dimensional latent vectors. AE, Denoising AE Sonobuoy signal data (simulated) MSE: 0.000238, 100x data reduction with ~4% reconstruction error. focussed on specific sonobuoy environments, limiting generalizability to other systems.
[88] Improved signal robustness under high antenna correlation with phase-shift keying and antenna index embedding. AE Frameworks with Antenna Index Embedding MIMO spatial modulation scenarios 18--24 dB gain in BLER and power efficiency (Rician factor of 20 dB). Limited testing under other channel conditions or configurations.
[90] Automated modulation feature extraction for rapid and high-accuracy classification. Stacked Denoising AE Synthetic modulation signals Superior accuracy and speed compared to conventional methods in noisy conditions. Lack of real-world signal testing and reliance on synthetic data.
[89] Enhanced SNR and reduced BER with improved feature extraction and reconstruction quality. DRdA-CA Eight modulation types (e.g., 16QAM, 64QAM) SNR improvement: 8.38 dB for 16QAM, BER reduction up to 63.5%. High computational complexity may limit real-time deployment.
[92] High classification accuracy under fading and noisy conditions with reduced reconstruction error. AE with non negativity Constraint Fourth-order cumulants of modulated signals 100% accuracy at 10-15 dB SNR, 86% at 0 dB SNR. Limited scalability to higher-dimensional or diverse datasets.
[94] Effective demodulation of high-frequency information with improved image quality. Hybrid Feature Modulated AE Channel modulated polarisation imaging data PSNR improvement: +180% (S3), SSIM: 0.9665. Limited exploration of diverse imaging scenarios beyond polarisation imaging.
[91] Enhanced denoising and wavelength demodulation for FSO links under turbulence. SDAE, DBN FSO communication signals with real noise SNR improvement: +22 dB, Wavelength accuracy: 0.06 nm. Dependency on specific FSO setups; generalisation to other noise types unverified.
[93] Efficient radio technology and modulation classification on low-cost platforms. LSTM Denoising AE RadioML2018.01A, Real-time PSD data 96.56% classification accuracy at high SNR, 4.3x faster classification speed on Raspberry Pi 4. Limited evaluation for advanced modulation schemes or other radio datasets.
X-A4 Cognitive Radio Networks (CRNs)
@@
Deep AEs improve CRN by improving spectrum sensing and interference mitigation, allowing efficient and adaptive communication in dynamic environments. For spectrum sensing, AEs extract essential features from signals, detect anomalies, and identify underused spectrum bands, with Denoising AEs proving effective in reconstructing clean signals under noisy conditions. In interference mitigation, AEs learn to suppress interference, adapt modulation schemes, and reconstruct signals in multi-user scenarios, which showed their ability to predict interference levels and dynamically adjust communication parameters. By automating feature extraction, providing real-time adaptability, and handling complex multidimensional data, AEs optimise spectrum use and ensure reliable, interference-aware communication in CRNs.
Spectrum Sensing and Anomaly Detection
Abdalzaher et al. [95] proposed a deep AE-based framework called (DAE-TRUST), to detect malicious nodes in cognitive radio (CR)-enabled IoT networks. By improving cooperative spectrum detection, the system achieved 100% accuracy in detecting malicious secondary users under specific conditions, with slight reductions in challenging environments. Subekti et al. [96] presented a spectrum sensing method that combines a deep AE for feature learning and an SVM for classification. This approach achieved up to 95% sensing accuracy in distinguishing primary and secondary users. Similarly, Nallarasan et al. [97] used AEs to detect random jamming attacks in cognitive radio IoT (CRIoT) networks, which achieved 89% detection precision by modelling jamming as an anomaly detection problem.
Signal Classification and Modulation Recognition
Liu et al. [98] introduced a stack hybrid AE framework for industrial CRNs, which achieved 97.75% classification accuracy in high SNR environments and thus outperformed other methods such as Capsule Networks. Zhu et al. [99] proposed using stacked denoising AEs for modulation classification, which achieved 97% accuracy at 5 dB SNR and surpassed traditional maximum likelihood methods in both classification accuracy and execution time efficiency.
Radio Map Estimation and Denoising
Teganya et al. [100] developed a deep-completion AE for estimating radio occupancy maps, which learnt spatial propagation patterns with fewer measurements compared to traditional methods. Almazrouei et al. [101] demonstrated the efficacy of convolutional denoising AEs in denoising radio signals, which achieved 96% accuracy at SNR = 0 dB and with robust performance under varying channel conditions.
Dynamic Interference Management and Beyond-5G
Wu et al. [102] proposed an AE based on adaptive deep learning (ADL) to manage dynamic interference in multi-user Gaussian channels, which significantly reduced BER and SER in strong interference scenarios. This approach enables adaptable constellations for 5G and beyond communication systems and demonstrated robustness in extreme interference conditions.
TABLE IX: Comparison of Studies on CRNs Using Deep AEs
@@
[96] Spectrum sensing for CR systems Deep AE + SVM Spectrum sensing data 95% sensing accuracy Limited to multi-radio environments
[97] Jamming attack detection AE CRIoT network data 89% detection accuracy Less effective for continuous jamming
@@
[99] Modulation classification Stacked Denoising AE Modulation data 97% accuracy at 5 dB SNR Higher complexity in dense symbol sets
@@
[102] Interference mitigation for beyond-5G systems Adaptive DL-based AE Multi-user Gaussian channel data Significant BER/SER improvement Higher complexity in real-time adaptation
TABLE X: Comparison of studies on deep AE-based methods for MIMO systems

### v2

@@ VI-A2 MIMO Systems
[97] flexible MC-AE for any channel coding Turbo-style Multi-Carrier AE OFDM Systems Achieves frequency diversity gain and surpasses OFDM-IM High computational cost due to multi-stage learning
VI-A3 OFDM Systems
Conventional OFDM systems face significant performance and efficiency limitations due to inherent waveform characteristics such as high peak-to-average power ratio (PAPR) and low adjacent channel leakage ratio (ACLR). High PAPR leads to nonlinear distortions in power amplifiers that degrade system performance [53, 35, 98]. Similarly, poor ACLR performance signifies inadequate spectral containment, which can result in increased interference with adjacent frequency bands and reduced spectral efficiency. Table VII lists the state-of-the-art comparison of AE-based models for OFDM systems.
The use of AE-driven waveform design methodologies that account for PAPR and ACLR constraints during training has been proposed [35, 49]. They focused on jointly optimizing a high-dimensional modulation scheme while satisfying the PAPR and ACLR constraints without compromising spectral containment or inducing power inefficiencies. The proposed AE used ResNet blocks with separable convolutional layers and the ReLU activation function, along with a separable two-dimensional (2D) convolution output layer. However, such approaches involve some limitations, such as difficulty in defining target values for PAPR and ACLR, complexity of computing true average energy, and limitation in accurately reflecting and optimizing for real-world channel conditions. In the same direction, Aoudia et al. [94] proposed a learning-based method for the joint design of transmit and receive filters, constellation geometry, and bit labeling, along with a neural network detector, evaluated on AWGN and 3GPP multipath channels. Their approach aimed to maximize the achievable information rate while satisfying ACLR and PAPR constraints.
Another critical aspect addressed by AE-based systems is synchronization in OFDM systems. Felix et al. [95] investigated the integration of an AE into an OFDM system with a CP. This was shown to mitigate synchronization issues and simplify equalization over multipath channels. Furthermore, their work demonstrated that the AE inherently learns to manage certain hardware impairments, such as non-linear amplifiers, without requiring explicit compensation algorithms.
Phase noise (PN) is another significant hardware impairment, which particularly impacts performance at higher frequencies like sub-THz as frequency and bandwidth increase. Coupled with the need for low PAPR in sub-THz due to amplifier limitations, constellation shaping becomes crucial for mitigating these issues. Marasinghe et al. [96] addressed the challenges of phase noise and PAPR in sub-THz communications through geometric constellation shaping. They proposed optimizing constellation shapes to be robust to PN while maintaining low PAPR under practical conditions, using an augmented Lagrangian approach for waveform optimization. Their results indicate a trade-off between PAPR and BLER performance, and stricter PAPR constraints can lead to reduced maximum amplitude and increased constellation point density. This work highlights the potential of tailored constellations to replace conventional QAM or APSK formats. The study mentions using filtered Gaussian techniques to generate phase noise samples.
Although these learning-based approaches show promise in PAPR/ACLR reduction, pilotless transmissions, and handle hardware impairments, several challenges and areas remain, such as complexity and generalization. The complexity of the used architecture, such as the use of ResNet, which often incurs high computational complexity in the training process. The generalization is that systems that are trained under a specific channel model often encounter performance deterioration when utilized in real-world channels, primarily due to the divergences that exist between the training model and the actual channel conditions.
Furthermore, Xu et al. [97] introduced a turbo-style multi-carrier AE (MC-AE) for OFDM systems. The MC-AE design allows for flexible DNN integration into channel coded systems and is capable of handling unknown channel models. The decoder within this framework is modified to produce reliable soft-bit decisions to achieve beneficial iteration gain through the exchange of extrinsic information with the channel decoder. Simulation results indicate substantial performance improvements of MC-AE compared to traditional methods, including conventional OFDM and OFDM-based index modulation in channel coded systems. However, there are limitations concerning the choice of channel coding schemes and parameters in deep learning integration. While the MC-AE demonstrates adaptability to unknown channel models, the interplay between the learned AE components and traditional channel coding elements requires careful consideration regarding parameter selection.
VI-A4 ISAC Systems
ISAC systems face critical challenges in jointly optimizing communication and sensing (JCAS) performance, primarily due to the inherently conflicting design requirements for each functionality. Communication systems typically prioritize constellations optimized for data throughput and robustness, whereas sensing applications demand constellations that enhance detection accuracy and resolution. This divergence results in a fundamental trade-off between the two objectives [99]. Table VII lists the state-of-the-art comparison of AE-based models for ISAC systems.
TABLE VIII: Comparison of Studies on AE-based ISAC systems
[99] Joint GSC and PSC for ISAC OFDM-ISAC, AWGN Simulated-based experiments Increase GMI by 0.16 bit/symbol while reducing kurtosis constraint by 0.19 hyperparameters fine-tuning, discrepancies between desired and achieved kurtosis levels
[100] Improved data transmission Dense (ReLU), symbol-wise Simulated, stochastic AWGN BLER below 10310{-3} at SNR >> 0 dB Evaluated only in AWGN
To balance these competing demands, constellation shaping has emerged as a pivotal strategy. However, achieving optimal performance requires careful consideration of both geometric and probabilistic constellation design, each of which offers distinct advantages under different operational constraints. Geometric shaping tends to be more effective under strict sensing constraints, while probabilistic shaping performs better when such constraints are relaxed. A notable challenge remains in optimizing detection probability while simultaneously satisfying performance and regulatory constraints, such as maintaining a fixed false alarm rate [99, 100]. Another approach to balance sensing and communication performance is the use of Joint constellation shaping using AE frameworks that has shown significant promise to enable simultaneous optimization of communication reliability and sensing accuracy [99]. These frameworks leverage novel loss functions that integrate both communication and sensing performance metrics to ensure balanced optimization. The discrepancy between analytical models and real-world sensing performance further complicates this optimization process, as a result it often leads to inconsistencies between simulated and empirical detection probabilities [99]. On the other hand, AE-based systems have been extended to support permutation-invariant encoding strategies, such as counting-based encoding schemes, which improve radar target detection in JCAS scenarios [101]. However, this permutation invariance of targets during detection poses training challenges for AEs. Therefore, modern ISAC systems, particularly those that implement JCAS, encounter difficulties in detecting multiple radar targets due to the inherent permutation invariance of the targets. This property introduces challenges for training AEs, as traditional loss functions and network architectures are not naturally suited for such scenarios [100, 101].
In RF signal classification, AEs have been used in the field of human sensing where AEs have been proposed to extract a compressed knowledge representation of the original input in RF sensing tasks [102]. On the other hand, Kompella et al. [103] proposed using a Vector-Quantized Variational Autoencoder (VQ-VAE) to augment training data for RF signal classification. They generate high-fidelity synthetic RF signals using the VQ-VAE. Experimental results demonstrate that augmenting datasets with this synthetic data significantly improves classification performance. However, the VQ-VAE equalizer requires additional computation, and it struggles to scale with high code dimensions.
VI-A5 CR Technology
@@
Deep AEs improve CRN by improving spectrum sensing and interference mitigation, allowing efficient and adaptive communication in dynamic environments. For spectrum sensing, AEs extract essential features from signals, detect anomalies, and identify underused spectrum bands, with Denoising AEs proving effective in reconstructing clean signals under noisy conditions. In interference mitigation, AEs learn to suppress interference, adapt modulation schemes, and reconstruct signals in multi-user scenarios, which showed their ability to predict interference levels and dynamically adjust communication parameters. By automating feature extraction, providing real-time adaptability, and handling complex multidimensional data, AEs optimize spectrum use and ensure reliable, interference-aware communication in CRNs. Table IX lists the state-of-the-art comparison of AE-based models for CR technology.
Abdalzaher et al. [104] proposed a deep AE-based framework called (DAE-TRUST) for detecting malicious nodes in CR-enabled IoT networks. The DAE-TRUST model employed a deep AE to identify anomalies in input data in ensuring spectrum access without collisions while protecting primary user rights by filtering out malicious reports. The effectiveness of the model was evaluated across six different environments, incorporating realistic conditions, including shadow fading measurements. A lack of prior research on AE-based trust models for security against jamming attacks in CR-assisted IoT networks puts this work as an early exploration in this specific area. However, challenges include the difficulties in integrating ML techniques into existing infrastructures without requiring additional hardware, the impact of large-scale fading on performance, and the complexity introduced by false reports affecting node reputation.
Liu et al. [105] introduced a stack hybrid AE framework for industrial CRNs. The proposed method is a semi-supervised classification approach that combines SHAE and a semi-supervised linear discriminant for effective feature extraction and classification. It effectively utilized CWD time-frequency analysis for feature extraction from wireless signals. The SHAE method outperformed AE, DAE, and SAE in classification accuracy, especially at low SNR. However, a significant challenge lies in the development of algorithms that generalize across various radio access scenarios and the need for solutions applicable across multiple radio standards.
Teganya et al. [106] developed a deep-completion AE for estimating radio occupancy maps, which learned spatial propagation patterns with fewer measurements compared to traditional methods. The approach addressed the challenge of accurately estimating radio maps due to complex electromagnetic wave propagation, which represents a significant area of interest in the field of cognitive radio networks. A key finding is that this method can significantly reduce the number of measurements required for accurate map estimation compared to existing schemes. However, this approach is sensitive to interference, based on the assumption that the measurements come from a single known transmitter. Finite-precision arithmetic can introduce significant errors in map estimates. Furthermore, a high computational complexity arises when predictions are needed in many locations.
Almazrouei et al. [107] demonstrated the efficacy of convolutional denoising AEs (CDAE) in denoising radio signals, which focused on denoising preambles from IEEE 802.11 protocols using simulated data. It achieved 96% accuracy at SNR = 0 dB and with robust performance under varying channel conditions. The use of CDAE for denoising radio signals is emphasized as a technique that can enhance classification performance within a machine learning pipeline. Critically, the complexity of applying denoising AE in a classification pipeline remains a challenge. In addition, there is a need for solutions that are applicable across multiple radio standards.
TABLE IX: Comparison of Studies on Deep Learning AEs for CR Technology
@@
VI-B Optical Fiber Communications
Optical fiber communication presents unique challenges, especially with increasing data rates and the need to overcome physical limitations like chromatic dispersion and nonlinearity, alongside complexities in signal processing such as equalization and carrier recovery in transceiver design. Figure 7 illustrates these identified challenges and also presents the solutions of AE-based methodologies.
Optical Fiber CommunicationsSolutionsChallenges Dispersion and Nonlinearity: CD, ISI, Kerr nonlinearity, long memory optical channel Equalization and Carrier Recovery: robust equalization and carrier recovery Constellation Shaping: non-Gray labeling, dedicated shaping encoder Joint design transmitter and receiver, RNN, Low-complexity offline learning, … active challenge area (VAE, VQ-VAE, VAE-LE) Equalizers, … active challenge area Joint GS and PS through AE, differentiable bit-level PS … active challenge area
Figure 7: Optical fiber communication challenges and solutions
VI-B1 Dispersion and Nonlinearity
Optical fiber communication systems face significant performance degradation due to chromatic dispersion and intersymbol interference (ISI) effects over fiber links [108, 109]. These impairments are particularly problematic in long-haul fiber scenarios, where achieving robustness against dispersion variations remains a critical challenge [110]. Furthermore, long memory of optical fibers complicates accurate modeling of interference effects across symbol sequences, which poses difficulties for both DSP techniques and end-to-end learning frameworks [108, 111]. Additional challenges emerge from the nonlinear behavior of the fiber, such as Kerr nonlinearity, which further restricts system capacity and complicates signal recovery [112].
To mitigate these impairments, Uhlemann et al. [112] explored the use of AE deep learning for the optimization of coherent optical communication systems, particularly under the influence of CD and Kerr nonlinearity. The fiber channel model was implemented using the split-step Fourier method (SSFM). The findings indicated that the AE could learn compensation and achieve high spectral efficiency even under challenging conditions. However, AE struggles to compensate for impairments at high input powers, and the training process may not consistently converge to a suitable solution. The interpretability of the AE learned behavior may also limit its performance. Furthermore, the lack of a closed-form solution for certain fiber models restricts the AE compensation capabilities, and training complexity remains a practical limitation, which requires careful adjustment of the AE structure.
AEs have facilitated the development of end-to-end optimization frameworks that jointly design transmitter and receiver components, bypassing the need for explicit mathematical models of complex fiber optic channels [109]. These frameworks have shown notable gains in performance and reach, especially when adapted to intensity modulation/direct detection (IMDD) systems. Recurrent neural network (RNN) architectures, including bidirectional and feedforward variants, have also been shown to be effective in capturing and compensating for channel memory effects, making them particularly suitable for dispersive optical environments [113, 25]. A differentiable link has been modeled that accounts for CD and memory effects in the optical fiber channel. The FFNN AE is noted as being unable to compensate for interference outside of the symbol block. In addition, diminishing returns in performance gains are observed as system nonlinearities and noise increase. However, high complexity and large model size of ANNs limit the effectiveness of AE architectures. Furthermore, such transceivers learned on specific model assumptions often perform poorly when applied to actual transmission links due to discrepancies, necessitating optimization of ANN parameters using experimental data. On the other side, finding the optimal symbol distribution for fiber channels remains an open problem, and the complexity of optimizing probabilistically shaped coherent links is a challenge. Furthermore, there is also a lack of exploration into the effects of various analog or digital impairments beyond those explicitly modeled.
In the same direction, low complexity offline learning approaches have demonstrated improved adaptability of NN-based transceivers to dispersion variations in highly dispersive IMDD links [110]. However, a primary difficulty lies in the need to develop efficient optimization procedures that remain robust across varying link conditions without relying on additional on-device learning, which can introduce prohibitive computational overhead.
Table X lists the state-of-the-art comparison of AE-based models for optical fiber communication.
TABLE X: Comparison of AE-based models in optical fiber communications